10) The initial section of the B chain of insulin is represented by the following amino acids: phenylalanine - valine - aspartic acid - glutamine - histidine - leucine - cysteine - glycine - serine - histidine. Determine the quantitative ratio of A + T and G + C in the DNA strand encoding this region of insulin.


1
Expert's answer
2021-02-17T12:40:50-0500

Nucleotide sequence (BLAST ncbi resourse)

TTC TAC ACA CCC AAG ACC CGC CGG GAG GCA 

AAG ATG TGT GGG TTC TGG GCG GCC CTC CGT (two strands)

percentage of A-T pairs = 11/30 = 36,7%

percentage of G-C pairs = 19/30 = 63,3%